OpenStudy (anonymous):

Find (f^-1)'(a): f(x) = 3+x^2+tan(pi(x)/2), -1

OpenStudy (anonymous):

\[(f^{-1})'(a)=\frac{ 1 }{f'(f ^{-1}(a)) }\]

OpenStudy (anonymous):

f(a); \[3=3+x ^{2}+\tan(\frac{ pix }{2 })\]

OpenStudy (anonymous):

\[0=x ^{2}+\tan(\frac{ pix }{ 2 })\]

OpenStudy (anonymous):

stuck

OpenStudy (anonymous):

If I am even on the right path.

OpenStudy (anonymous):

I'm a little confused how you took the derivative.

OpenStudy (anonymous):

Actually one sec.

OpenStudy (anonymous):

Did you first take the inverse and then the derivative?

OpenStudy (anonymous):

Have not done the derivative yet: doing: \[(f ^{-1})(x) = y <-> f(y) = x\]

OpenStudy (anonymous):

so if I place 3 in for my f(x), I should find out what my \[(f ^{-1})\] is correct?

OpenStudy (anonymous):

After taking the derivative.

OpenStudy (anonymous):

so f'(x) = 2x+sec^2(pix/2)?

OpenStudy (anonymous):

poop did I do an integral there with the sec^2?

OpenStudy (anonymous):

\[f'(x) = 2x+\frac{ 1 }{ 2 }\pi[\sec ^{2}(\frac{ x \pi }{ 2 })]\]

OpenStudy (anonymous):

then ..... 3 = the lovely right side above?

OpenStudy (anonymous):

@jim_thompson5910 If you have time.

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

these inverses are going to be the death of me.

OpenStudy (anonymous):

Was either path an ok way to go? Me initially and then what z suggested.

jimthompson5910 (jim_thompson5910):

hint: what is f(0)

OpenStudy (anonymous):

f(0) = 3

jimthompson5910 (jim_thompson5910):

take the inverse of both sides f(0) = 3 f^-1[f(0)] = f^-1(3) 0 = f^-1(3) f^-1(3) = 0

jimthompson5910 (jim_thompson5910):

do you see what to do now?

OpenStudy (anonymous):

man my head is spinning.

OpenStudy (anonymous):

not sure. put 0 into the f'(x) and solve?

OpenStudy (anonymous):

If I did the derivative correct.

jimthompson5910 (jim_thompson5910):

You are correct (f^-1)'(a) = 1/[f ' ( f^-1(a) )] (f^-1)'(3) = 1/[f ' ( f^-1(3) )] (f^-1)'(3) = 1/[f ' ( 0 )] (f^-1)'(3) = ???

OpenStudy (anonymous):

\[f'(3) = 2(3)+\frac{ 1 }{ 2 }\pi[\sec ^{2}(\frac{ 3 \pi }{ 2 })]\]

OpenStudy (anonymous):

or am i missing yet another simple step and am killing myself with the algebra?

jimthompson5910 (jim_thompson5910):

you want f'(0), not f'(3)

OpenStudy (anonymous):

\[0+\frac{ 1 }{ 2 }\pi(1)\] \[=\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

suppose i should do a couple hundred and then it may start to get easier.

OpenStudy (anonymous):

I miss college algebra.

OpenStudy (anonymous):

Thank you both for your help!

jimthompson5910 (jim_thompson5910):

np